OpenStudy (anonymous):

find dy/dx of y=x^7(3^x)

OpenStudy (anonymous):

\[(fg)'=f'g+fg'\]

OpenStudy (anonymous):

dy/dx = f'(x)g(x) + f(x)g(x) = (x^7)' * (3^x) + (x^7) * (3^x)' = [7x^6 * 3^x] + [x^7 * ln(3) * 3^x] = 3^x * [7x^6 + [x^7 *ln(3)]]

OpenStudy (anonymous):

good but \[(3^x)'=x \ln 3*3^x\]

OpenStudy (anonymous):

Hmmm. x*ln(3) would imply there that the natural log was ln(3^x), but that's not the case if you look at the original definition: d/dx (a^x) = ln(a) * a^x Check it out: http://www.math.com/tables/derivatives/more/b%5Ex.htm

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2

OpenStudy (anonymous):

what about xy=cot^3(y+x)^2